# Vulnerability Management Chaining: An Integrated Framework for Efficient Cybersecurity Risk Prioritization

### v2

@@ 6.4 Limitations and Future Research Directions
An important consideration for this research is the rapidly evolving vulnerability management landscape. New predictive metrics such as Likely Exploited Vulnerabilities (LEV) and enhanced versions of existing systems (e.g., EPSS v4.0) have emerged during our study period, while security vendors have independently developed integration approaches conceptually similar to our methodology. These developments underscore both the timeliness of our research question and the industry-wide recognition that single-metric vulnerability management approaches are fundamentally insufficient for modern threat environments.
@@
Several factors may limit the generalizability of our specific findings. Our exploitation datasets, while comprehensive, represent specific organizational contexts and time periods. Performance may vary in different threat environments or against different attack patterns. Vulnerability management performance may change over time as threat actor behaviors evolve, new exploitation techniques emerge, or defensive capabilities improve. Different organizations face different threat profiles based on their industry, size, geographic location, and security posture.
@@
Extended studies tracking framework performance over multiple years could identify temporal patterns and inform adaptive management strategies. Research examining how different types of organizations adapt and customize the framework could inform best practices and implementation guidance. Studies examining how framework performance changes as threat landscapes evolve could inform proactive adaptation strategies. Improved methods for identifying and validating vulnerability exploitation could strengthen evaluation methodologies for all vulnerability management research.
Organizations considering adoption of our framework should consider gradual deployment with specific vulnerability classes or systems before full-scale deployment to build confidence and identify necessary customizations. Systematic tracking of framework performance against organizational metrics including time to remediation, exploitation incidents, and resource utilization can inform continuous improvement efforts. Staff training on threat intelligence interpretation and framework customization supports successful implementation.
The demonstrated effectiveness of KEV integration supports continued investment in government-led threat intelligence sharing initiatives. Professional organizations and standards bodies should consider incorporating threat intelligence integration into vulnerability management guidance and frameworks. Regulators requiring vulnerability management programs should consider allowing evidence-based prioritization approaches as alternatives to pure CVSS-based requirements.
@@ 7 Conclusion
This paper presents Vulnerability Management Chaining, a novel framework that integrates CVSS, EPSS, and KEV to achieve more efficient vulnerability prioritization while maintaining comprehensive coverage. Our research demonstrates that systematic integration of multiple intelligence sources can overcome the fundamental limitations that plague individual vulnerability management approaches.
7.1 Key Contributions and Practical Impact
This paper presents Vulnerability Management Chaining, a novel methodological framework that integrates CVSS, EPSS, and KEV to achieve more efficient vulnerability prioritization while maintaining comprehensive coverage. Our research demonstrates that systematic integration of multiple data sources can overcome the fundamental limitations that plague individual vulnerability management approaches.
@@ 7.1 Key Contributions and Practical Impact
Our analysis reveals that KEV and EPSS identify complementary sets of exploited vulnerabilities, with their combination capturing 57 additional vulnerabilities that neither method would identify individually. By relying exclusively on freely available data sources, our framework enables broad adoption without requiring expensive commercial threat intelligence subscriptions.
@@ 7.2 Future Directions and Final Remarks
As the cybersecurity field continues to evolve toward more sophisticated, data-driven approaches, frameworks like ours will become increasingly important for managing complex security decisions at scale. To support reproducibility and enable further research, we commit to making our experimental data, analysis code, and framework implementation available through open source repositories. This comprehensive availability aligns with our goal of making advanced vulnerability management accessible to organizations regardless of their resources or expertise levels.

### v3

@@ 6.3 Evolution of Threat Intelligence Landscape
These developments underscore both the timeliness of our research question and the industry-wide recognition that single-metric vulnerability management approaches are fundamentally insufficient for modern threat environments.
@@
Several factors may limit the generalizability of our specific findings. Our exploitation dataset, while comprehensive, represents specific organizational contexts and time periods. Performance may vary in different threat environments or against different attack patterns. Vulnerability management performance may change over time as threat actor behaviors evolve, new exploitation techniques emerge, or defensive capabilities improve. Different organizations face different threat profiles based on their industry, size, geographic location, and security posture.
@@
Extended studies tracking framework performance over multiple years could identify temporal patterns and inform adaptive management strategies. Research examining how different types of organizations adapt and customize the framework could inform best practices and implementation guidance. Investigation of how emerging threat intelligence sources could be incorporated while maintaining the framework’s simplicity and effectiveness would ensure continued relevance as the threat intelligence ecosystem evolves.
@@ 7 Conclusion
This research addresses a critical challenge in modern cybersecurity: the overwhelming volume of vulnerabilities that organizations must evaluate and prioritize for remediation. Through systematic integration of complementary vulnerability intelligence sources, we demonstrate that dramatic efficiency improvements are achievable while maintaining comprehensive security coverage.
7.1 Summary of Contributions
Our Vulnerability Management Chaining framework represents the first systematic methodology for integrating CVSS technical severity assessment, EPSS predictive modeling, and KEV confirmed exploitation evidence into a unified decision framework. This integration addresses fundamental limitations in each individual approach while preserving their unique strengths. Our research demonstrates that systematic integration of multiple data sources can overcome the fundamental limitations that plague individual vulnerability management approaches.
@@ 7.1 Summary of Contributions
Our analysis reveals that KEV and EPSS identify complementary sets of exploited vulnerabilities, with their combination capturing 48 additional vulnerabilities that neither method would identify individually. By relying exclusively on freely available data sources, our framework enables broad adoption without requiring expensive commercial threat intelligence subscriptions.
@@ 7.2 Future Directions and Final Remarks
We encourage security practitioners to implement and adapt our framework to their specific organizational contexts, and we invite researchers to build upon our methodology to develop even more effective approaches to this critical challenge. Only through continued innovation in vulnerability management can defenders hope to keep pace with the ever-expanding threat landscape.